OpenStudy (brainzonly):

The most challenging mathematical puzzle of all time: Archimedes' Riddle... If thou art diligent and wise, O stranger, compute the number of cattle of the Sun, who once upon a time grazed on the fields of the Thrinacian isle of Sicily...

OpenStudy (brainzonly):

Complete text of the riddle: PART I If thou art diligent and wise, O stranger, compute the number of cattle of the Sun, who once upon a time grazed on the fields of the Thrinacian isle of Sicily, divided into four herds of different colors, one milk white, another glossy black, the third yellow and the last dappled. In each herd were bulls, mighty in number according to these proportions: Understand, stranger, that the white bulls were equal to a half and a third of the black together with the whole of the yellow, while the black were equal to the fourth part of the dappled and a fifth, together with, once more, the whole of the yellow. Observe further that the remaining bulls, the dappled, were equal to a sixth part of the white and a seventh, together with all the yellow. These were the proportions of the cows: The white were precisely equal to the third part and a fourth of the whole herd of the black; while the black were equal to the fourth part once more of the dappled and with it a fifth part, when all, including the bulls, went to pasture together. Now the dappled in four parts were equal in number to a fifth part and a sixth of the yellow herd. Finally the yellow were in number equal to a sixth part and seventh of the white herd. If thou canst accurately tell, O stranger, the number of Cattle of the Sun, giving separately the number of well-fed bulls and again the number of females according to each color, thou wouldst not be called unskilled or ignorant of numbers, but not yet shalt thou be numbered among the wise. PART II But come, understand also all these conditions regarding the cattle of the Sun. When the white bulls mingled their number with the black, they stood firm, equal in depth and breadth, and the plains of Thrinacia, stretching far in all ways, were filled with their multitude. Again, when the yellow and the dappled bulls were gathered into one herd they stood in such a manner that their number, beginning from one, grew slowly greater till it completed a triangular figure, there being no bulls of other colours in their midst nor none of them lacking. If thou art able, O stranger, to find out all these things and gather them together in your mind, giving all the relations, thou shalt depart crowned with glory and knowing that thou hast been adjudged perfect in this species of wisdom.

OpenStudy (brainzonly):

Have fun and see you in 100 years! LOL

OpenStudy (brainzonly):

this is in Diophantine analysis....

OpenStudy (brainzonly):

@Seratul Can you figure it out? :D I know it's \[7.76*10^{something}\] it stumped thousands of people this is so exciting I love problems that have practically never been solved. it's supposed to be a trick: it's got 2 answers conceived 2 different ways but only 1 of those is correct...

OpenStudy (seratul):

That's too much to read :'(

OpenStudy (brainzonly):

LOL!!!!!

OpenStudy (brainzonly):

I've ALMOST got it....

OpenStudy (brainzonly):

White bulls=\[\left( \frac{ 1 }{ 2 }+ \frac{ 1 }{ 3 }\right)\]black bulls+yellow bulls

OpenStudy (brainzonly):

Black Bulls= \[\left( \frac{ 1 }{ 4 }+\frac{ 1 }{ 5 } \right)\] dappled bulls +yellow bulls

OpenStudy (brainzonly):

Dappled Bulls=\[\left( \frac{ 1 }{ 6 }+\frac{ 1 }{ 7 } \right)\] white +yellow bulls

OpenStudy (brainzonly):

White Cows=\[\left( \frac{ 1}{ 3 } +\frac{ 1 }{ 4 }\right)\] black herd

OpenStudy (brainzonly):

Black Cows=\[\left( \frac{ 1 }{ 4 } +\frac{ 1 }{ 5 }\right)\] dappled herd

OpenStudy (brainzonly):

Dappled Cows=\[\left( \frac{ 1 }{ 5 }+\frac{ 1 }{ 6 } \right)\] yellow herd

OpenStudy (brainzonly):

Yellow Cows=\[\left( \frac{ 1 }{ 6 }+\frac{ 1 }{ 7 } \right)\] white herd

OpenStudy (brainzonly):

yay that's over next step..

OpenStudy (brainzonly):

White bulls + black bulls = a square number, Dappled bulls + yellow bulls = a triangular number.

OpenStudy (brainzonly):

let's write the number of white, black, dappled, and yellow bulls as W,B,D, and Y and the number of white, black, dappled, and yellow cows as w , b , d , and y...

OpenStudy (brainzonly):

???it creates a system of seven equations with eight unknowns!!! :(

OpenStudy (brainzonly):

50,389,082 is the total but it's not the answer! :D

OpenStudy (brainzonly):

the solution given above for the first part of the problem should be multiplied by \[n=\frac{ \left( w ^{4658j}-w ^{-4658j} \right)^{2} }{ \left( 4657 \right)\left( 79072 \right) }\]

OpenStudy (brainzonly):

where\[w=300426607914281173363\sqrt{609} + 84129507677858393258\sqrt{7766}\] and j is any positive integer

OpenStudy (brainzonly):

Equivalently, squaring w makes: \[w ^{2}=u+v \sqrt{(609)(7766)}\] where (u, v) are fundamental solutions of the Pell Equation (which takes forever to explain, pardon me :D) which is \[u ^{2}-(609)(766)v ^{2}=1\]

OpenStudy (brainzonly):

If we put 1 in the place of j, this will make the smallest herd satisfying both equations. Let's skip this part....

OpenStudy (brainzonly):

*and satisfies both parts of the problem..

OpenStudy (brainzonly):

So the size of the smallest herd is..... \[7.76*10^{206544}\] There you go! :) tell me if I'm wrong :D

OpenStudy (brainzonly):

Don't try solving the exponent, you'll get an "error" on your calculator and the number won't fit in your house, much less on your tiny computer screen!!!

OpenStudy (legomyego180):

the answer is 42

OpenStudy (brainzonly):

???????????????????????????????????????????????????????????????????????????????????????????????????????????????????????????? 42 COWS???????

OpenStudy (legomyego180):

The life, universe and everything does include cows last time I checked.

OpenStudy (brainzonly):

there's no way its 42. Any calculator will tell u it's a number too big. u try it!!!

OpenStudy (markchernioglo):

Too big with decimal, or too big whole number?

OpenStudy (markchernioglo):

You could probably round if its a decimal number

OpenStudy (brainzonly):

10^206544= INVALID INPUT./thank you very much.

OpenStudy (brainzonly):

and no its an exponent.

OpenStudy (markchernioglo):

oh just saw it. Geez thats a huge number

OpenStudy (brainzonly):

they've only calculated it once and i'm not even going to TRY posting it on here because you'd be reading #s for the rest of your life.

OpenStudy (markchernioglo):

Haha I bet! I mean thats a huge exponent!

OpenStudy (brainzonly):

yeah and then multiply 7.76 to it!!!!! have fun. totally! I mean it. XD

OpenStudy (brainzonly):

I think all my brilliant calculations are slowing down the page's loading time LOL

OpenStudy (markchernioglo):

Nope not fun... :p

OpenStudy (brainzonly):

^lol

OpenStudy (brainzonly):

thanks for the medal bro I need those :D

OpenStudy (markchernioglo):

haha same

OpenStudy (markchernioglo):

lol this problem is soooo confusing.. It is too huge!

OpenStudy (brainzonly):

lol I took 8 days solving it bfor I posted it on here

OpenStudy (markchernioglo):

lol yea it would probably take me 8 years...

OpenStudy (brainzonly):

Check it out: If thou art able, O stranger, to find out all these things and gather them together in your mind, giving all the relations, thou shalt depart crowned with glory and knowing that thou hast been adjudged perfect in this species of wisdom I am perfect???? O.O And crowned with glory??? :O

OpenStudy (brainzonly):

give me a minute, I have to pick my jaw up off the floor XD

OpenStudy (markchernioglo):

Lol bro you are a comedian haha

OpenStudy (brainzonly):

Thanks. everybody needs a good joke every now and then :D

OpenStudy (markchernioglo):

Haha especially one dealing with math. :D

OpenStudy (brainzonly):

LOL